• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name seely

    Log entry time 10:21:09 on November29,2006

    Entry number 189468


    Flow from CHL dropped to 0 for unknown reasons (they are looking at
    possible vacuum problems). Once supply was reopened target condensed.
    Cryo requests that we keep flow high for now. JT3 is at 50%, HPH at
    650W. CHL flow 10g/s.ad to change gas bottle and add hydrogen to
    system. Now at ~20psi at 19K. Reset some low pressure and high power
    alarm limits.